OpenStudy (anonymous):

suppose 8000 is invested at interest rate k, compounded continuously, and grows to 10962.07 in 5 yr. Find the interest rate. HELLLLPPPP

OpenStudy (anonymous):

Using this formula A=P(e)^kt

OpenStudy (anonymous):

substitute into this formula: \[A=p(1+\frac{ r }{ n })^{nt}\]

OpenStudy (anonymous):

so your p= 8000, a= 10962.07, t=5

OpenStudy (anonymous):

ok ill try that

OpenStudy (anonymous):

sorry the formula is \[A=pe ^{rt}\]

OpenStudy (anonymous):

just substitute and to get the rate out of the exponent do ln on both sides of the equation

OpenStudy (anonymous):

dangit i dont know why but im stiil having trouble sorry.. im horrible at word problems

OpenStudy (anonymous):

\[10962.07=8000e ^{r5}\]

OpenStudy (anonymous):

\[\frac{ 10962.07 }{ 8000 }=e ^{r5}\]

OpenStudy (anonymous):

\[137.025875=e ^{r5}\]

OpenStudy (anonymous):

i mean i got to that but when i get 1.3703is that the answer or how do i go about the e ^r5

OpenStudy (anonymous):

ln both sides

OpenStudy (anonymous):

\[\ln 1.37025875=\ln e ^{r5}\]

OpenStudy (anonymous):

that will allow to get the r5 out of the exponent: \[lne=1\]

OpenStudy (anonymous):

so do you multiply the 5 by the 1.37025875

OpenStudy (anonymous):

divide so that you can isolate r

OpenStudy (anonymous):

i got like .27405175

OpenStudy (anonymous):

I messed up somewhere, sorry , let me see

OpenStudy (anonymous):

okay lol

OpenStudy (anonymous):

It should give us the interest rate !

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

are you sure you have the numbers right?. I plug in the wrong numbers all the time. lol

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

I got it: 6.3%

OpenStudy (anonymous):

r=6.3%

OpenStudy (anonymous):

since r=0.6299991812 multiplied by 1oo since it's a percentage and rounded out

OpenStudy (anonymous):

\[0.06299991812x1000=6.299 \]

OpenStudy (anonymous):

6.299=6.3%

OpenStudy (anonymous):

if you plug in r in the original equation, it will give you 10962.07, but you have to put that whole crazy number.

OpenStudy (anonymous):

OHHHH OK i see

OpenStudy (anonymous):

funny, because I have a test tomorrow with these same kind of formulas.

OpenStudy (anonymous):

really, i have a test next week

OpenStudy (anonymous):

well I hope I helped you out. I learned too at the same time.

OpenStudy (anonymous):

haha yea you did thank you. i appreciate your help

OpenStudy (anonymous):

sure no problem. I have been helped here by other people too. take care